OpenStudy (anonymous):

If f (x) = −x 2 − 2x + 5, find f (−3)

OpenStudy (anonymous):

You just need to put -3 in as x.

OpenStudy (anonymous):

oh sweet!!! ok!

OpenStudy (anonymous):

f(-3) = - (-3)2 -2(-3) + 5 = - 9 + 6 + 5 = 2.